OpenStudy (usukidoll):

@satellite73 need your help figuring this out.... going to type in details in the comments

OpenStudy (usukidoll):

I had email my prof and he gave me three hints

OpenStudy (anonymous):

oh no not again i was confused enough the first time

OpenStudy (usukidoll):

yeah but I've email my prof and he gave me these hints and stuff... would that help a bit? O_O

OpenStudy (anonymous):

you tax my mind but i guess i can try but first i have to say i have no idea what \(f_{\infty}\) means

OpenStudy (anonymous):

we got @zarkon there too, maybe he can help

OpenStudy (usukidoll):

1.) Show that \(f_\infty \)is defined, i.e., if \(a \in A_\infty\), then \(f(a) \in B_\infty.\) 2.) Show that \(f_\infty\) is injective, clear because \(f_\infty\) is a restriction of the injective function f. 3.) Show that \(f_\infty \)is onto, i.e., if\( b \in B_\infty\), then there is an \(a \in A_\infty\) and \(f(a) = b\). So, show that there is an appropriate element in A that maps to b, and then show that it even belong to \(X_\infty\).

OpenStudy (usukidoll):

and each of those 3 hints have 2 line proofs... which I don't get hth is that possible

OpenStudy (usukidoll):

AHHHHH! DJPON3!!! she's like my fav pony!

OpenStudy (anonymous):

the only thing i can think of is that \(A_{\infty}\) is an infinite sequence of elements of \(A\) and that \(f_{\infty}\) is the extension of \(f\) so that \(f(\{a\}_n)=\{b\}_n\) where \(b_i=f(a_i)\)

OpenStudy (usukidoll):

ah we could try that approach! At least that's something

OpenStudy (anonymous):

what i am trying to say is that i don't know what \(f_{\infty}\) means

OpenStudy (usukidoll):

:/ me either... but how the heck did my classmates score a 9 out of 10 on this thing?

OpenStudy (anonymous):

probably knew what it meant!

OpenStudy (anonymous):

tag @Zarkon maybe he knows what it means

OpenStudy (usukidoll):

nugh I ham so stuck!!!!

OpenStudy (anonymous):

do you have somewhere a definition of \(f_{\infty}\)? without it, there is no point in thinking about it

OpenStudy (usukidoll):

I could look it up ...

OpenStudy (anonymous):

I can't help unless I know what \(f_\infty\) is.

OpenStudy (usukidoll):

found this line from the link Consider the sets Ainf and Binf. Every b in Binf obviously has an inverse under f in Ainf. Therefore f is a bijection from Ainf to Binf.

OpenStudy (usukidoll):

f infinity the set that doesn't terminate if it has a sequence idk better ask again. x.x

OpenStudy (usukidoll):

there sent an email ... going to drop this for now...until tomorrow.

OpenStudy (usukidoll):

@satellite73 @ wio .. I found out what \[f_\infty\] is.. it's the Restriction of f, with domain and codomain restricted

OpenStudy (usukidoll):

@wio